Hello, Can someone please review my answer to the following question. Please help me.
A person spins the pointer and is awarded the amount indicated by the pointer. Circle with left half $2.00 the upper right quarter is $12 and the lower right quarter is $5.00 and it cost $6 to play the game. There are only 3 part to the circle


It costs $6 to play the game. Determine:
These are my answers I am not sure they are right
• The expectation of a person who plays the game.
E(x) = (1/2)2 + (1/4)12 + (1/4)5
E = 1+3+1.25 =5.25

•My Answer:The fair price to play the game. Wheel
Fair price = 6/3 = 2.00

Answer by stanbon(75887)

A person spins the pointer and is awarded the amount indicated by the pointer. Circle with left half $2.00 the upper right quarter is $12 and the lower right quarter is $5.00 and it cost $6 to play the game. There are only 3 part to the circle

It costs $6 to play the game. Determine:
These are my answers I am not sure they are right
• The expectation of a person who plays the game.
E(x) = (1/2)2 + (1/4)12 + (1/4)5
E = 1+3+1.25 =5.25
----
Your expected winnings are 5.25. But you
paid 6 to play the game. So your expected
"winnings" are really -0.75. You can expect
to lose 75 cents each time you play the game.
==============================================
To make the game "fair", the vendor should only
charge $5.25 to play the game.
=================================
Cheers,
Stan H.